40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 27 16:26:11 2019, Yehonathan, Update, PSL, PMC Linewidth measurement PMCTransmissionSpectra.pdf
    Reply  Wed Jan 1 14:14:00 2020, Yehonathan, Update, PSL, PMC Linewidth measurement 150MHzSideBandCreation.pdfLinewidthMeasurment.pdf
       Reply  Wed Jan 15 15:04:31 2020, gautam, Update, PSL, PMC Linewidth measurement 
          Reply  Fri Jan 17 22:11:19 2020, Yehonathan, Update, PSL, Ringdown measurements 20200117_192455.jpg20200117_192448.jpg
             Reply  Sat Jan 25 11:54:42 2020, Yehonathan, Update, PSL, Ringdown measurements Input_beam_path.jpgReflected_Beam_Path.jpgTransmitted_Beam_Path.jpgPMCRingdownNormalizedRawdata.pdfTransPDFits.pdf
                Reply  Mon Jan 27 21:35:06 2020, Yehonathan, Update, PSL, Ringdown measurements IMC_REFL_Beam_Path.jpg
                   Reply  Mon Jan 27 21:48:49 2020, gautam, Update, PSL, Ringdown measurements elevatedSeis.pdf
                      Reply  Tue Jan 28 16:01:17 2020, gautam, Update, IOO, IMC WFS servos stable again QPD_recenter.png
                         Reply  Tue Jan 28 20:51:37 2020, Yehonathan, Update, IOO, IMC WFS servos stable again 
                            Reply  Wed Jan 29 12:40:24 2020, Yehonathan, Update, IOO, IMC Ringdowns preliminary data analysis IMCRingdownNormalizedRawdata.pdfIMCTransPDFits.pdf
                               Reply  Mon Feb 3 13:54:10 2020, Yehonathan, Update, IOO, IMC Ringdowns extended data analysis IMCTransReflAnalysis_anotated.pdf
                                  Reply  Wed Feb 5 21:13:17 2020, Yehonathan, Update, IOO, IMC Ringdowns extended data analysis 
                                     Reply  Sat Feb 15 12:56:24 2020, Yehonathan, Update, IOO, IMC Transfer function measurement IMCFrequencyResponse.pdfIMCRingdownNormalizedRawdata.pdfIMCREFLPDFits.pdf
Message ID: 15132     Entry time: Fri Jan 17 22:11:19 2020     In reply to: 15126     Reply to this: 15154
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: Ringdown measurements 

I prepare for the ringdown measurement of the PMC according to Gautam's previous experiments.

1. I assembled the needed PDs and power supplies, lenses, beamsplitters and optomechanics needed for the measurement.

2. I surveyed the laser power with an Ophir power meter in the different parts of the experiment. All the measurements were done with the AOM driver excited with 1V DC.

For the PMC reflection, we chose to split off the beam that goes into the reflection camera. The power in that beam is ~ 0.11mW when the PMC is locked and 2.1mW otherwise.

For the PMC transmission, we chose to split the beam that is transmitted through the second steering mirror after the PMC. The power in that beam is 2mW.

For the peak off before the PMC, we chose to split the beam that goes into the fiber coupler. That path contains also the other AOM diffraction orders: 2.26mW in the 0th order beam, 6.5mW in the 1st order beam, 0.14mW in the 2nd order beam.

3. I placed a 10% beam splitter in the peak-off path such that 90% still goes into the fiber coupler (Attachment 1). I place a lens and PDA255 to measure the peak-off (Attachment 2).

It's getting late, I'll continue with the PD placements on Tuesday.

Attachment 1: 20200117_192455.jpg  843 kB  | Show | Hide all | Show all
Attachment 2: 20200117_192448.jpg  1.367 MB  | Hide | Hide all | Show all
20200117_192448.jpg
ELOG V3.1.3-